# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v1

@@ 5.1.3 $G_{25}$
The group $G_{25}$ of order 648 is also denoted $W(L_{3})$ with degrees $\deg(G_{25})=\{6,9,12\}$ [18, Table D.3]. $G_{25}$ has two conjugacy classes, consisting of matrices with nontrivial eigenvalue $w$ or $w^{2}$ respectively, for $w=\zeta_{3}$. Expanding the base field to be the cyclotomic field of order 36, we get 8 possible types, which actually split into 4 pairs of inverses. We only list one set of inverse eigenvalues for each pair in Table 5.
@@ 5.1.4 $G_{26}$
The group $G_{26}$ of order 1296 is also denoted $W(M_{3})$ with degrees $\deg(G_{26})=\{6,12,18\}$ [18, Table D.3]. In $G_{26}$ there are three conjugacy classes of reflections, corresponding to nontrivial eigenvalues in $\{\zeta_{3},\zeta_{3}^{2},-1\}$. In this group the are 6 sets of eigenvalues for the inverse product, which correspond to 3 inverse pairs, so we only list one from each pair.
@@ 5.3.1 $G_{28}$
The group $G_{28}$ of order 1152 is also denoted $W(F_{4})$ with degrees $\deg(G_{28})=\{2,6,8,12\}$ [18, Table D.3]. All reflections in $G_{28}$ have nontrivial eigenvalue -1, and form two conjugacy classes of reflections. We only found one type of nice 4-tuples in this group, with two possible choices of $\lambda$ parameters for middle convolution.
@@ 5.3.2 $G_{30}$
The group $G_{30}$ of order 14400 is also denoted $W(H_{4})$ with degrees $\deg(G_{30})=\{2,12,20,30\}$ [18, Table D.3]. All reflections in $G_{30}$ have nontrivial eigenvalue -1 and are conjugate to each other. Looking at the degrees, we extend the base field to $\mathbb{Q}(\zeta_{60})$. We found three distinct types of nice 4-tuples, the details are in Table 21.
@@ 5.3.3 $G_{32}$
The group $G_{32}$ of order 155520 is also denoted $W(L_{4})$ [18, Table D.3]. In this group, the nontrivial eigenvalues of reflections are in $\{\zeta_{3},\zeta_{3}^{2}\}$ corresponding to two conjugacy classes. For $G_{32}$, $\deg(G_{32})=\{12,18,24,30\}$ so the eigenvalues of multiplicity 2 must be 12-th roots of unity. Expanding the field to $\mathbb{Q}(\zeta_{12})$ we found 5 distinct eigenvalue sets for the inverse product. Note that to ensure we found all examples, we would have to run the code over $\mathbb{Q}(\zeta_{360})$ since there may be matrices with one or two eigenvalues other than the 12-th roots which would not be found, however we claim it is enough to check over $\mathbb{Q}(\zeta_{12})$.

### v2

@@ 5.1.3 $G_{25}$
The group $G_{25}$ of order 648 is also denoted $W(L_{3})$ with degrees $\deg(G_{25})=\{6,9,12\}$ [21, Table D.3]. $G_{25}$ has two conjugacy classes, consisting of matrices with nontrivial eigenvalue $w$ or $w^{2}$ respectively, for $w=\zeta_{3}$. Expanding the base field to be the cyclotomic field of order 36, we get 8 possible types, which actually split into 4 pairs of inverses. We only list one set of inverse eigenvalues for each pair in Table 5.
@@ 5.1.4 $G_{26}$
The group $G_{26}$ of order 1296 is also denoted $W(M_{3})$ with degrees $\deg(G_{26})=\{6,12,18\}$ [21, Table D.3]. In $G_{26}$ there are three conjugacy classes of reflections, corresponding to nontrivial eigenvalues in $\{\zeta_{3},\zeta_{3}^{2},-1\}$. In this group the are 6 sets of eigenvalues for the inverse product, which correspond to 3 inverse pairs, so we only list one from each pair.
@@ 5.3.1 $G_{28}$
The group $G_{28}$ of order 1152 is also denoted $W(F_{4})$ with degrees $\deg(G_{28})=\{2,6,8,12\}$ [21, Table D.3]. All reflections in $G_{28}$ have nontrivial eigenvalue -1, and form two conjugacy classes of reflections. We only found one type of nice 4-tuples in this group, with two possible choices of $\lambda$ parameters for middle convolution.
@@ 5.3.2 $G_{30}$
The group $G_{30}$ of order 14400 is also denoted $W(H_{4})$ with degrees $\deg(G_{30})=\{2,12,20,30\}$ [21, Table D.3]. All reflections in $G_{30}$ have nontrivial eigenvalue -1 and are conjugate to each other. Looking at the degrees, we extend the base field to $\mathbb{Q}(\zeta_{60})$. We found three distinct types of nice 4-tuples, the details are in Table 21.
@@ 5.3.3 $G_{32}$
The group $G_{32}$ of order 155520 is also denoted $W(L_{4})$ [21, Table D.3]. In this group, the nontrivial eigenvalues of reflections are in $\{\zeta_{3},\zeta_{3}^{2}\}$ corresponding to two conjugacy classes. For $G_{32}$, $\deg(G_{32})=\{12,18,24,30\}$ so the eigenvalues of multiplicity 2 must be 12-th roots of unity. Expanding the field to $\mathbb{Q}(\zeta_{12})$ we found 5 distinct eigenvalue sets for the inverse product. Note that to ensure we found all examples, we would have to run the code over $\mathbb{Q}(\zeta_{360})$ since there may be matrices with one or two eigenvalues other than the 12-th roots which would not be found, however we claim it is enough to check over $\mathbb{Q}(\zeta_{12})$.